# Constraint Preconditioning and Parameter Selection for a First-Order Primal-Dual Method applied to Model Predictive Control

### v1

@@ I INTRODUCTION
Modified Ruiz equilibriation [6] and the hypersphere preconditioner [7] are two existing preconditioning techniques for first-order primal-dual algorithms. These preconditioners aim to reduce the condition number of the KKT matrix. Both perform a change of variables and additionally, modified Ruiz equilibriation scales the constraints with a diagonal matrix. These two preconditioner transform set $\mathbb{D}$ in Problem I since they both employ a change of variables. This requires special structure in the matrices defining the change of variables in order to preserve closed-form projections onto $\mathbb{D}$, which is needed to solve Problem I by using Proportional-Integral Projected Gradient (pipg), a first-order primal-dual optimization algorithm [8, 9, 10].
@@ II Proportional-Integral Projected Gradient
Proportional-Integral Projected Gradient (pipg), shown in Algorithm 1, is a first-order primal-dual conic optimization algorithm capable of infeasibility detection. This algorithm is specialized to handle common constraints sets which arise in optimal control problems, such as boxes, balls, and second-order cones, via closed-form projections [8, 9, 10]. PIPG has been demonstrated to be faster than many state-of-the art optimization solvers such as ecos, mosek, gurobi, and osqp [10]. pipg has been applied extensively to trajectory optimization problems including three-degree-of-freedom (DoF) powered-descent guidance and has been used as a subproblem solver within sequential convex programming for multi-phase powered-descent guidance, 6-DoF powered-descent-guidance, and spacecraft rendezvous problems [7, 11, 12, 13, 14].
@@ III-A Derivation
First, we derive a preconditioner that only modifies Constraint 1b of Problem I and minimizes the condition number of the KKT matrix. First, we need the following assumptions on the constraint matrix $H$ and the Hessian of the objective function $P$.
@@
The KKT conditions of Problem I are as follows
$\underbrace{\begin{bmatrix}P&H^{\top}\\ H&0_{m\times m}\end{bmatrix}}_{\mathcal{K}}\begin{bmatrix}x^{\star}\\ w^{\star}\end{bmatrix}+\begin{bmatrix}q\\ -g\end{bmatrix}\in\begin{bmatrix}-\mathbb{N}_{\mathbb{D}}(z^{\star})\\ \{0_{m}\}\end{bmatrix}$ (3)
@@
where $H^{\top}=QR$ is the thin or economy QR factorization of $H^{\top}$, $\hat{H}=\eta Q^{\top}$, $\hat{g}=\eta R^{-\top}g$, and Assumption 1 ensures invertibility of $R^{\top}$ [16]. Note that Assumption 1 is equivalent to requiring that linear independence constraint qualification (LICQ) of Constraint 1b holds for Problem I [17]. We can arbitrarily scale the constraint in Equation 5d by some $\eta>0$ and the rows of $\hat{H}$ will still be orthogonal. If we choose $\eta=\sqrt{\lambda_{\mathrm{max}}\lambda_{\mathrm{min}}+\lambda_{\mathrm{min}}^{2}}$, then $\hat{H}\hat{H}^{\top}=(\lambda_{\mathrm{min}}\lambda_{\mathrm{max}}+\lambda_{\mathrm{min}}^{2})I_{m}$, thus minimizing the condition number of the KKT matrix. The QR preconditioner is given by Algorithm 2.
@@ III-B Geometric Interpretation
We can see that in Figure 1, pipg takes several iterations to converge. Notice that Line 2 of pipg takes a step in the negative gradient direction of the objective function in order to minimize the objective, then takes a step in a direction given by a linear combination of the rows of $H$ (the $H^{\top}w^{k+1}$ term) to satisfy Constraint 1b. The rows of $H$ are the normal vectors of the hyperplanes which define the equality constraints in Problem I. If we have two equality constraints with nearly parallel hyperplane normals (as in Figure 1), we require smaller coefficients in $w^{k+1}$ for the linear combination of the hyperplane normals, given by $H^{\top}w^{k+1}$, to move the iterates in a direction nearly parallel to hyperplane normals (up and to the right in Figure 1), but require large coefficients for the iterates to move in a direction nearly orthogonal to the hyperplane normals (down and to the right in Figure 1). The coefficients of this linear combination are given by our current iterate of the dual variables $w^{k+1}$. To get large enough $w^{k+1}$ for the iterates to move nearly orthogonal to the hyperplane normals, we must wait for the integrator given by Line 1 to accumulate enough error, resulting in slow convergence.
@@ IV PARAMETER SELECTION
There are infinite choices for $\alpha$ and $\beta$ which satisfy the above relationship, and it has been observed that for fastest convergence a “proper” choice for $\alpha$ and $\beta$ must be made. From Proposition 1 of [8], we see that zero primal-dual gap is a sufficient condition for optimality of Problem I. Thus, we propose to choose step-sizes which minimize the upper bound on the primal-dual gap.
@@
We rewrite Theorem 1 from [8], which provides an upper bound on the primal-dual gap for Algorithm 1 to explicitly show dependence on $\gamma$
@@
where $\{z^{j},w^{j}\}_{j=1}^{k}$ are generated by Algorithm 1, $(z^{\star},w^{\star})$ is the optimal primal-dual solution of Problem I, and $\mathcal{L}(z,w)=\frac{1}{2}z^{\top}Pz+q^{\top}z+w^{\top}(Hz-g)$ is the Lagrangian of Problem I. Both $\gamma$ and $\sigma$ are defined to be positive real numbers. Under this assumption, we see that $f(\gamma)$ given by Equation 8 is convex in $\gamma$ and has a minimizer given by
@@ V-D Analysis of results
where $z^{\star}$ is the solution to I computed by Gurobi [19] using YALMIP [20]. We terminate pipg when the relative error, $\mathrm{error}_{\mathrm{opt}}$, decreases below $1.0\times 10^{-4}$, indicating the solution has a relative error of $0.01\%$.

### v2

@@ I INTRODUCTION
Modified Ruiz equilibriation [6] and the hypersphere preconditioner [7] are two existing preconditioning techniques for first-order primal-dual algorithms. These preconditioners aim to reduce the condition number of the KKT matrix. Both perform a change of variables and additionally, modified Ruiz equilibriation scales the constraints with a diagonal matrix. These two preconditioner transform set $\mathbb{D}$ in Problem 1 since they both employ a change of variables. This requires special structure in the matrices defining the change of variables in order to preserve closed-form projections onto $\mathbb{D}$, which is needed to solve Problem 1 by using Proportional-Integral Projected Gradient (pipg), a first-order primal-dual optimization algorithm [8, 9, 10].
@@ II Proportional-Integral Projected Gradient
Proportional-Integral Projected Gradient (pipg), shown in Algorithm 1, is a first-order primal-dual conic optimization algorithm capable of infeasibility detection. This algorithm is specialized to handle common constraints sets which arise in optimal control problems, such as boxes, balls, and second-order cones, via closed-form projections [8, 9, 10]. pipg has been demonstrated to be faster than many state-of-the art optimization solvers such as ecos, mosek, gurobi, and osqp [10]. It has also been extensively applied to trajectory optimization problems including three-degree-of-freedom (DoF) powered-descent guidance and has been used as a subproblem solver within sequential convex programming for multi-phase powered-descent guidance, 6-DoF powered-descent-guidance, and spacecraft rendezvous problems [7, 11, 12, 13, 14].
@@ III-A Derivation
First, we derive a preconditioner that only modifies Constraint 1b of Problem 1 and minimizes the condition number of the KKT matrix. First, we need the following assumptions on the constraint matrix $H$ and the Hessian of the objective function $P$.
@@
The KKT conditions of Problem 1 are as follows
$\underbrace{\begin{bmatrix}P&H^{\top}\\ H&0_{m\times m}\end{bmatrix}}_{\mathcal{K}}\begin{bmatrix}z^{\star}\\ w^{\star}\end{bmatrix}+\begin{bmatrix}q\\ -g\end{bmatrix}\in\begin{bmatrix}-\mathbb{N}_{\mathbb{D}}(z^{\star})\\ \{0_{m}\}\end{bmatrix}$ (3)
@@
where $H^{\top}=QR$ is the thin or economy QR factorization of $H^{\top}$, $\hat{H}=\eta Q^{\top}$, $\hat{g}=\eta R^{-\top}g$, and Assumption 1 ensures invertibility of $R^{\top}$ [16]. Note that Assumption 1 is equivalent to requiring that linear independence constraint qualification (LICQ) of Constraint 1b holds for Problem 1 [17]. We can arbitrarily scale the constraint in Equation 5d by some $\eta>0$ and the rows of $\hat{H}$ will still be orthogonal. If we choose $\eta=\sqrt{\lambda_{\mathrm{max}}\lambda_{\mathrm{min}}+\lambda_{\mathrm{min}}^{2}}$, then $\hat{H}\hat{H}^{\top}=(\lambda_{\mathrm{min}}\lambda_{\mathrm{max}}+\lambda_{\mathrm{min}}^{2})I_{m}$, thus minimizing the condition number of the KKT matrix. The QR preconditioner is given by Algorithm 2.
@@ III-B Geometric Interpretation
We can see that in Figure 1, pipg takes several iterations to converge. Notice that Line 2 of pipg takes a step in the negative gradient direction of the objective function in order to minimize the objective, then takes a step in a direction given by a linear combination of the rows of $H$ (the $H^{\top}w^{k+1}$ term) to satisfy Constraint 1b. The rows of $H$ are the normal vectors of the hyperplanes which define the equality constraints in Problem 1. If we have two equality constraints with nearly parallel hyperplane normals (as in Figure 1), we require smaller coefficients in $w^{k+1}$ for the linear combination of the hyperplane normals, given by $H^{\top}w^{k+1}$, to move the iterates in a direction nearly parallel to hyperplane normals (up and to the right in Figure 1), but require large coefficients for the iterates to move in a direction nearly orthogonal to the hyperplane normals (down and to the right in Figure 1). The coefficients of this linear combination are given by our current iterate of the dual variables $w^{k+1}$. To get large enough $w^{k+1}$ for the iterates to move nearly orthogonal to the hyperplane normals, we must wait for the integrator given by Line 1 to accumulate enough error, resulting in slow convergence.
@@ IV PARAMETER SELECTION
There are infinite choices for $\alpha$ and $\beta$ which satisfy the above relationship, and it has been observed that for fastest convergence a “proper” choice for $\alpha$ and $\beta$ must be made. From Proposition 1 of [8], we see that zero primal-dual gap is a sufficient condition for optimality of Problem 1. Thus, we propose to choose step-sizes which minimize the upper bound on the primal-dual gap.
@@
Theorem 1 from [8] provides an upper bound on the primal-dual gap for Algorithm 1. We will rewrite this theorem to explicitly show dependence on $\gamma$
@@
where $\{z^{j},w^{j}\}_{j=1}^{k}$ are generated by Algorithm 1, $(z^{\star},w^{\star})$ is the optimal primal-dual solution of Problem 1, and $\mathcal{L}(z,w)=\frac{1}{2}z^{\top}Pz+q^{\top}z+w^{\top}(Hz-g)$ is the Lagrangian of Problem 1. Both $\gamma$ and $\sigma$ are defined to be positive real numbers. Under this assumption, we see that $f(\gamma)$ given by Equation 8 is convex in $\gamma$ and has a minimizer given by
@@ V-D Analysis of results
where $z^{\star}$ is the solution to 1 computed by Gurobi [19] using YALMIP [20]. We terminate pipg when the relative error, $\mathrm{error}_{\mathrm{opt}}$, decreases below $1.0\times 10^{-4}$, indicating the solution has a relative error of $0.01\%$.